Assume function is as follows:

 

           C =   500 + 0.75Yd

                          

a) Determine the national income if investment is 200 using expenditure and injection leakage approach. (5marks)

 

b) Draw the national income equilibrium (consumption and saving curves) based on information in question (a). (5marks)

 

a)     The following are some assumptions for an economy:

·        Marginal propensity to consume (MPC) is 0.75

·        Autonomous consumption is $6,000 million.

·        Autonomous investment is $3,500 million.

                                                                

What is the level of equilibrium national income using expenditure and injection-leakage approach? (5marks)



1
Expert's answer
2021-01-29T20:00:33-0500

а) The national income will be:

Yd = C + I

Yd = 500 + 0.75Yd + 200

Yd = 700 + 0.75Yd

0.25Yd = 700

Yd = "\frac{700}{0.25}=2800"

The national income will be 2800.

b) Consumption curve will be:



The saving curve:


In graph 1 black lines shows the consumption function corresponding to each level of income whereas 45-degree line represents income. Consumption curve intersects the income curve at poit where income = consumption, that is why this point is the break-even point which shows zero saving. It states that saving curve in graph 2 must intersect the x-axis at this level of income. Also, to the left of this break-even point, consumption is more than income which depicts dissaving or negative saving while to the right of this break-even point income is more than consumption, which shows that there is positive saving, the same is depicted in graph 2.

In graph 1 the slope pf consumption function is 0.75, this is the marginal propensity to consume (MPC), while in graph 2, the slope is 0.25, which is the marginal propensity to save (MPS), as MPC + MPS = 1. MPC was already given in the consumption function as 0.75.

c) MPC is 0.75

Autonomous consumption is 6000 million and autonomous investment is 3500 million.

The level of national income will be:

C = 6000 + 0.75Yd

I = 3500

Yd = C +I = 6000 + 0.75Yd + 3500 = 9500 + 0.75Yd

0.25Yd = 9500

Yd = "\frac{9500}{0.25} = 38000"

The national income which will achieve equilibrium would be 38000 million.